Use (15.26) to calculate Djj for O2 at 0°C and (a) 1.00 atm; (b) 10.0 atm. The experimental value at 0°C and 1 atm is 0.19 cm2/s.
1
Expert's answer
2020-05-04T01:52:35-0400

v = u / p

a) 0.19 / 1 = 0.19 Djj

b) 0.19 / 10 = 0.019 Djj